I'm going to assume you're looking for the probability of getting three heads out of three coin spins and that you're using a fair coin. For coin spins, theoretical probability is very simple. The probability of getting three heads in a row is 1/2 * 1/2 * 1/2 = 1/8. This means that if you tossed a coin three times, you'd expect to see three heads once every 8 trials. For experimental probability you need to define clear trials, for this experiment you can't just spin a coin over and over and count the number of times you see three heads in a row, for example, if you threw the following: H T H H T T H H H H H T T H T T T you have three cases where you have three heads in a row, but they all overlap so these are not independent trials and cannot be compared to the theoretical result. When conducting your experiment, you know that if you get a T in your trial, it doesn't matter what comes after, that trial has already failed to get three heads in a row. The trial is deemed a success if you get three heads in a row, naturally. As a result, if you threw the above sequence, you would to determine your experimental probability in the following way: H T fail H H T fail T fail H H H success H H T fail T fail H T fail T fail T fail In this example we have 8 trials and one success, therefore the experimental probability is 1/8. The sample variance (look it up), however is also 1/8, meaning that all you really know is that the experimental probability could be anywhere between 0 and 1/4. The only way to get the variance down (and therefore reduce your confidence interval) is to perform more and more trials. It's unlikely for the theoretical probability and experimental probability to be EXACTLY the same but the more trials you do, the more the experimental probability will converge on the theoretical probability.
Draw a circle. Everyone puts marbles in the circle and tries to shoot them out of the circle . When you shoot one out it becomes yours. A larger marble is used as the "shooter".
What is the probability of getting a black face card?
There are 6 black face card in a deck of 52 cards, so the probability of getting a black face card is 6/52 = 3/26
Denote by mu the mean of X. Given
P(X>5.52)=0.305
i.e.,
P((X-mu)/4> (5.52-mu)/4)=0.305
So, letting Z=(X-mu)/4. Then Z~N(0,1). Hence, the above equality is
P(Z>(5.52-mu)/4)=0.305. So,
(5.52-mu)/4=0.51
So,
mu=5.52-4*0.51=3.48
So, the best answer is 3.5
A card is drawn from a pack of 52 cardsWhat is the probability that the card is an ace or a king?
513
What does it mean when an escort saids she is a spinner?
"Spinner" (Adjective - Slang) Used among some men to describe an extremely petite young woman, usually with a lot of sex appeal - the term is thought to reference the small stature and weight of the woman; (IE she is so petite she could be spun around while riding the man during sex). Ex: Man #1- "Aw man, look at that girl, she's so tiny." Man #2- "yeah, she's a spinner."
What is the probability of a coin?
If you mean what is the probability of getting a heads/tails, it is a 1 in 2 chance (50/50 chance). You are just as likely to get a heads as you are to get a tails.
1/4
What is the shape of this distribution?
A company is studying the number of daily debit card purchases. If there were 20 purchases and the probability of a debit card purchase is 0.5
What is the shape of this distribution
A three digit number can be chosen in 9*10*10 = 900ways
No. of ways to choose a three digit number without 7 is 8*9*9 = 648 ways
probability of picking a three digit number that includes atleast one digit that is 7 is 1- (648/900) = 252/900 = 7/25
Why is sample size important in determining probability?
The bigger the sample size the more accurate the results will be.
For example, if you roll a 6 sided die and track the results to get the probability of rolling a six. If you only roll 6 times, then you may not even get one 6 or you could get a few. A small sample size means you won't get very reliable results.
idk i wish that some would tell me lol COOKIE all day HHS HASS PROGRAM
all that matters in this problem, is that 2 of the 13 spaces are orange.
On the first spin, the odds are 11:2 against Jake getting the orange space.
On the second spin, the result of the first spin has no relevance or influence at all, so the odds are still 11:2 against.
Why is this question flagging up in "breakups"?
What is the probability of a Hole in one?
Average golfer making hole in one on designated hole: 1 in 12,500
Professional golfer making hole in one on designated hole: 1 in 2,500 Here's a link http://www.prlog.org/10092942-hole-in-one-insurance-odds-golfs-rare-feat.html
What is the probability of getting all heads or all tails if you flip a coin three times?
The probability of getting all heads if you flip a coin three times is:
P(HHH) = 1/2 ∙ 1/2 ∙ 1/2 = 1/8.
The probability of getting all tails if you flip a coin three times is:
P(TTT) = 1/2 ∙ 1/2 ∙ 1/2 = 1/8.
The probability of getting all heads or all tails if you flip a coin three times is:
P(HHH or TTT) = P(HHH) + P(TTT) = 2/8 = 1/4.
0 (there is no chance it will be short since tall is dominant over short).
Hope this helps!
- Biology Student
What is the product rule and the sum rule of probability?
Sum Rule:
P(A) = \sum_{B} P(A,B)
Product Rule:
P(A , B) = P(A) P(B|A) or P(A, B)=P(B) P(A|B)
[P(A|B) means probability of A given that B has occurred]
P(A, B) = P(A) P(B) , if A and B are independent events.
What is the probability that a child will have a round face brown or blond hair?
Depends on the parents.
The probability of getting heads twice and tails once if a coin is tossed 3 times?
3/8, or 37.5%
There are eight possible sequences that can emerge from three coin tosses. Three have heads twice and tails once. Three have the opposite, two tails and one heads. The remaining possibilities are all heads or all tails.
Why can't the probability of an event occurring be greater than 1?
If probability becomes greater than one, then the uncertainty coefficient fluctuates and results in a terminal case.